OpenStudy (anonymous):

How do you get the answer to this problem : A company earns a weekly profit of P dollars by selling x items, according to the equation P(x)=-5x^2+40x-300. How many items does the company have to sell each week to maximize the profit?

OpenStudy (zzr0ck3r):

find the max, and then find that x value. do you know calculas?

OpenStudy (anonymous):

No I don't ... I'm just in algebra 1

OpenStudy (zzr0ck3r):

what about how to find a vertex?

OpenStudy (anonymous):

I got 30 and 10 first. Then tried again and got -4

OpenStudy (zzr0ck3r):

-b/(2a) will give the vertex, the y value then use that to solve for x

OpenStudy (zzr0ck3r):

-5x^2+40x-300, so -40/(2(-5)) = 4

OpenStudy (anonymous):

I got 4 for -b divided by 2 times negative 5

OpenStudy (anonymous):

YeA. Got that. Thanks But how do u find the maximum

OpenStudy (zzr0ck3r):

this is your x value, now plug that into the function to get your y value

OpenStudy (zzr0ck3r):

that y value is the max

OpenStudy (anonymous):

But the answer is a negative number..... -220

OpenStudy (anonymous):

Would -220 be the max?

OpenStudy (anonymous):

O.. Wait I think I understand now 1 sec

OpenStudy (zzr0ck3r):

yes that is your max it does not make much since to me in context but thats it

OpenStudy (anonymous):

Yea same here. Thanks. :) Also can u please help me with this problem : the highs h in ft of a projectile launched vertically upward from the tip of a 96 foot tall tower when the time is given by t=0 is given by h=96+80t-16t^2. The object is launched when the time is 0. How LNG will it take the projectile to strike the ground? What is the maximum height that the projectile reaches?

OpenStudy (zzr0ck3r):

set it = to -96 and solve to get the time it will take to hit the broung, and then do the same thing we did last time to get the height but this time add 96 feet

OpenStudy (zzr0ck3r):

broung = ground

OpenStudy (zzr0ck3r):

does this make since?

OpenStudy (anonymous):

Not really. I am still trying to get the first part of the Answer

OpenStudy (zzr0ck3r):

now that i look at the function it looks like it has taken into acound the 96 feet, so just set the function equal to zero and solve. do you know how to do this>?

OpenStudy (anonymous):

I got 5 And -16 for the first part. Is that correct?

OpenStudy (anonymous):

Yes when it's equal to zero I got -1 and 6

OpenStudy (zzr0ck3r):

so 6 (units of time) is the answer to the first part

OpenStudy (anonymous):

Oh ok thanks. So to fine the next part do I do what I did for the first problem

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

it will be 196 feet

OpenStudy (anonymous):

Wait I don't understand how u got that. Sorry

OpenStudy (anonymous):

I got 160

OpenStudy (zzr0ck3r):

ok so the vertex is -80/(2(-16)) = 5/2 when I plug this into the function I get 196

OpenStudy (zzr0ck3r):

make since?

OpenStudy (anonymous):

Yea. Thanks. So to graph that would I just...

OpenStudy (zzr0ck3r):

you could plot all the points from 0-6 or use software to do it http://www.wolframalpha.com/input/?i=graph+96%2B80t-16t%5E2

OpenStudy (zzr0ck3r):

yeah?

OpenStudy (anonymous):

O ok. And Icahn use that for the first problem too. To graph?

OpenStudy (anonymous):

*** I can

OpenStudy (zzr0ck3r):

sure just type it in like I did at the top

OpenStudy (anonymous):

Thanks so much. You were a great help.

OpenStudy (zzr0ck3r):

np you can msg me anytime for help

OpenStudy (anonymous):

I need help with the geometry proof. > Please Given: angle AXY is congruent to the angle AYX, and angle AXC is congruent to angle AYB Prove: segment BY is congruent to segment CX

OpenStudy (anonymous):

please help me with this. thank you :) if you would like me to send the pic. just let me know